OpenStudy (anonymous):

Algebra 1 question! please help! Will post graph in a bit

OpenStudy (anonymous):

stock A B C Price f(x) = 25(1.08)x x f(x) The initial price of this stock is $30, but it has 0 22.00 been increasing 4% each day 1 21.56 2 21.13

OpenStudy (anonymous):

Create the price function for Stock D. It should have the second lowest starting price and the second highest rate of growth. Using complete sentences, justify that your function meets these requirements. Graph the price function for Stock D. Explain the key features of the graph with complete sentences. A recent drop in sales has affected Stock D with the function g(x) = –6. Explain to Gordon how Stock D’s new price function, f(x) + g(x), will be created. Graph f(x) + g(x). Gordon wants to fully understand what kind of changes could affect the money he is investing. Using complete sentences, explain the effect of the following transformations. Graph them and f(x) on the same coordinate plane. f(x + 2) f(–x) f(x) + 3

Mehek (mehek14):

flvs algebra 1?

OpenStudy (anonymous):

stock A B C Price f(x) = 25(1.08)x x f(x) The initial price of this stock is $30, 0 22.00 but it has been increasing 4% each day 1 21.56 2 21.13

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

@kidrah69

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

@Jamierox4ev3r

OpenStudy (anonymous):

i think i have an answer for the first part

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do you still need help?

OpenStudy (anonymous):

yes i do :)

OpenStudy (anonymous):

alright so...

OpenStudy (anonymous):

A recent drop in sales has affected Stock D with the function g(x) = –6. Explain to Gordon how Stock D’s new price function, f(x) + g(x), will be created. Graph f(x) + g(x). Gordon wants to fully understand what kind of changes could affect the money he is investing. Using complete sentences, explain the effect of the following transformations. Graph them and f(x) on the same coordinate plane. f(x + 2) f(–x) f(x) + 3

OpenStudy (anonymous):

i need help with the last 2 parts

OpenStudy (anonymous):

wait do you need help on the first post?

OpenStudy (anonymous):

you mean this post?

OpenStudy (anonymous):

oh sorry ok

OpenStudy (anonymous):

what was the function you created

OpenStudy (anonymous):

which means it started at 23 dollars and has a growth rate of 5% right?

OpenStudy (anonymous):

thats right

OpenStudy (anonymous):

"A recent drop in sales has affected Stock D with the function g(x) = –6. Explain to Gordon how Stock D’s new price function, f(x) + g(x), will be created. Graph f(x) + g(x)."

OpenStudy (anonymous):

this means f(x)+g(x) OR 23(1.05)^x-6

OpenStudy (anonymous):

ahh i see...

OpenStudy (anonymous):

Are you familiar which the graphing software GeoGebra?

OpenStudy (anonymous):

not really, i can look it up and figure it out

OpenStudy (anonymous):

https://www.geogebra.org/

OpenStudy (anonymous):

downloading it now

OpenStudy (anonymous):

press start creating and you need to log in with your google account

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Once it finishes downloading, you need to input the function 23(1.05)^x-6

OpenStudy (anonymous):

i added it to chrome now what?

OpenStudy (anonymous):

essentially your graph should look something like this

OpenStudy (anonymous):

i see

OpenStudy (anonymous):

so in your solution for "A recent drop in sales has affected Stock D with the function g(x) = –6. Explain to Gordon how Stock D’s new price function, f(x) + g(x), will be created. Graph f(x) + g(x)." you can write.... f(x)+g(x) = 23(1.05)^x-6 (and post the graph under it)

OpenStudy (anonymous):

i understand, now whats next?

OpenStudy (anonymous):

For the last question...Gordon wants to fully understand what kind of changes could affect the money he is investing. Using complete sentences, explain the effect of the following transformations. Graph them and f(x) on the same coordinate plane.

OpenStudy (anonymous):

this is basically just substituting

OpenStudy (anonymous):

f(x) stands for the function of x

OpenStudy (anonymous):

the exponent (^x) that you see at the end of your function is the same thing

OpenStudy (anonymous):

For example, your first one is f(x+2) right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

By substituting, your new function would be f(x+2)23(1.05)^(x+2)

OpenStudy (anonymous):

are you familiar which f(1)23(1.05)^1 f(2)23(1.05)^2 f(3)23(1.05)^3 f(4)23(1.05)^4

OpenStudy (anonymous):

these are the same things

OpenStudy (anonymous):

ok so just plugging in the x values

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

f(x+2)=23(1.05)^(x+2) f(-x)=23(1.05)^(-x) f(x)+3=23(1.05)^x+3

OpenStudy (anonymous):

these would essentially be your substituted functions

OpenStudy (anonymous):

you just need to put them into a graphing software and apply all of them on one graph

OpenStudy (anonymous):

sorry my computer froze

OpenStudy (anonymous):

It's ok

OpenStudy (anonymous):

this picture is the graph for all your functions

OpenStudy (anonymous):

y = f(x) where y is the dependent variable and x is the independent variable. This means when the independent variables changes, the dependent variable also changes. The function can shift and the money he is investing will either drop or increase. For example, f(x) + 3 means that the y-intercept will go up by three and Gordon’s stock will increase.

OpenStudy (anonymous):

Do you understand? any questions I can answer?

OpenStudy (anonymous):

i dont know how to graph them...

OpenStudy (anonymous):

oops take away the f(x+2) from the function, remember it is already inserted at the end of the function, like instead of f(x+2)=23(1.05)^(x+2), put in 23(1.05)^(x+2)

OpenStudy (anonymous):

ok i did it and my graph is still blank

OpenStudy (anonymous):

want me to take a picture?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

The manage bar is on stop of GeoGebra, the last icon looks like a compass rose

OpenStudy (anonymous):

press on it and press "move graph views"

OpenStudy (anonymous):

This way you can drag the graph, by dragging it to the bottom right a little bit, you should see your results

OpenStudy (anonymous):

23(1.3)^(x+2) 23(1.3)^(-x) 23(1.3)^x+3

OpenStudy (anonymous):

Put all these in on the same graph and it is the solution

OpenStudy (anonymous):

ohh ok thanks man!

OpenStudy (anonymous):

to explain the last question, y = f(x) where y is the dependent variable and x is the independent variable. This means when the independent variables changes, the dependent variable also changes. The function can shift and the money he is investing will either drop or increase. For example, f(x) + 3 means that the y-intercept will go up by three and Gordon’s stock will increase.

OpenStudy (anonymous):

So he can lose or gain money depending on the function xD

OpenStudy (anonymous):

lol ok, thanks man really appreciate all the help! gave you a medal and fan

OpenStudy (anonymous):

Alright you're welcome, peace!

OpenStudy (anonymous):

peace :)